A gas absorbs a photon of 355 nm and emits at two wavelengths. If one of the emissions is at
680 nm, the other is at ??

Grade:11

1 Answers

Avinash
8 years ago
Logic:

Since energy is conserved, the energy, Eabsof absorbed photon is equal to total of energies, (E1+ E2) of two photons emitted.

i.e. Eabs= E1+ E2

Where:

E1= Energy of first photon emitted

E2= Energy of second photon emitted

Energy, E and wavelength, λ of a photon are related as follows:

E = hc/λ

Where:

h = Planck's constant

c = velocity of light (a constant)

Therefore:

1/λabs= 1/λ1+ 1/λ2